12=2.43*W^0.3326: 12=2.43*W^0.3326
Answer by nabla(409)

12=2.43*w^0.3326
1200=243w^0.3326
1200/243=w^.3326
(1200/243)^(1/.3326)=w
121.70617262...=w

check:
2.43(121.70617262)^0.3326 is approximately= 12.